OpenStudy (misssunshinexxoxo):

Solve -2x + 1 less than or greater to -7 and describe the graph of the solution.

OpenStudy (misssunshinexxoxo):

I am having trouble please explain step by step

OpenStudy (misssunshinexxoxo):

@tester97

OpenStudy (istim):

HAve you went to purplemath.com before?

OpenStudy (misssunshinexxoxo):

no yet khan academy

OpenStudy (misssunshinexxoxo):

I usually don't go to sites like this unless I have to

OpenStudy (tester97):

\[-2x+1\le-7\] Is this how it looks?^^^

OpenStudy (misssunshinexxoxo):

Inqualities have been hard for me

OpenStudy (misssunshinexxoxo):

yes

OpenStudy (tester97):

ok so subract 1 from both sides what do you get|?

OpenStudy (misssunshinexxoxo):

1?

OpenStudy (misssunshinexxoxo):

Please show me step by step then I should understand it

OpenStudy (tester97):

yes subtract 1 from both sides because you are trying to find x right?

OpenStudy (misssunshinexxoxo):

Yes

OpenStudy (tester97):

so what do you get when you subtract 1 from both sidez?

OpenStudy (misssunshinexxoxo):

-1?

OpenStudy (misssunshinexxoxo):

@hero

OpenStudy (tester97):

no it would look like \[-2x \le-8\]

OpenStudy (tester97):

then u divide by neg 2 so that means u have to flip the sign and it would look like this \[x \ge 4\]

OpenStudy (misssunshinexxoxo):

how did you arrive to negative 2? where is it in the equation? I am still not getting this? Please help